OpenStudy (anonymous):

Find the slope of the line that contains the following points. A(5, 6), B(10, 8) 5/2 2/5 14/15

OpenStudy (anonymous):

dy/dx = (8-6)/(10-5) = 2/5

OpenStudy (anonymous):

thanks ill give you a medal